The amount of Principal is $3000, and the annual interest rate is 12%. Answer the following questions.

(a) What is the effective annual interest rate if compounded monthly? Calculate Total payback after 2 years in a lump sum.

(b) What is the simple annual interest rate to generate the same amount of payback as (a) in two years with the principal of $3000?
Answer #1

a) effective annual interest rate = (1 + nominal interest rate/12)12 - 1

effective annual interest rate = 1.0112 - 1 = 12.6825%

total payback = 3000 x (1 + 12.6825%)2 = $ 3809.2

b)

Payback = $ 3809.2

time = 2 years

simple annual rate = r

3000 + 3000 x r x 2/100 = 3809.2

r = 13.487%